OpenStudy (awkwardpanda):

Rewrite to make it easy to graph using a translation. Describe the graph?

OpenStudy (awkwardpanda):

\[^3\sqrt{27x-81}-5\]

OpenStudy (shamil98):

is the 5 inside the sqrt symbol?

OpenStudy (awkwardpanda):

nope :3

OpenStudy (shamil98):

\[\huge \sqrt[n]{x^m} = x^\frac{ m }{ n }\] \[\huge \sqrt[3]{27x-81} -5 =(27x-81)^{\frac{1 }{ 3 }} - 5\]

OpenStudy (awkwardpanda):

thanks dr. shum.

OpenStudy (shamil98):

mhm